Integrate : x^2+3/(x^6[x^2+1])
Please help

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
You can do the following:
\int \frac{x^2+3}{x^6\left ( x^2+1 \right )} dx=\int \frac{1}{x^6} dx+\int \frac{2}{x^6\left ( x^2+1 \right )} dx
Now using the partial fraction, we get:
\frac{2}{x^6\left ( x^2+1 \right )} =\frac{A}{x^2+1}+\frac{B}{x^2}+\frac{C}{x^4}+\frac{D}{x^6}
Solving for the constants, would give the required result.
Regards